OpenStudy (anonymous):

how do you figure these out?: 1. P(8,2) 2. C(12,4) * C(8,3) 3. C(5,2) 4. C(12,7)

OpenStudy (amistre64):

the are nPr and nCr

OpenStudy (amistre64):

you simply need to do them on a calculator for the most part for ease

OpenStudy (amistre64):

P(8,2) = 8*7 = 56

OpenStudy (anonymous):

yes! what she said! im doing this online school and they dont teach youanything!

OpenStudy (amistre64):

C(5,2) i ssimple enough as well; 1,5,10,10,5,1; id go with 5

OpenStudy (amistre64):

C(n,r) is a shorter version of: P(n,r)/r! the "!" might be off cause i cant recall it clearly

OpenStudy (amistre64):

C(5,3) should be 10 P(5,3)/3! = 5.4.3/3.2 = 5.2 = 10 ... id say the factorial notation is good

OpenStudy (amistre64):

C(12,7) 12.11.10.9.8.7.6 ---------------- 7.6.5.4.3.2 6.11.2.3.2 = 66.12 = 12(60+6) = 720+72 = 792? gonna have to dbl chk me steps

OpenStudy (amistre64):

yay its right

OpenStudy (amistre64):

C(12,4) * C(8,3) 12.11.10.9 8.7.6 ---------- ----- 4.3.2 3.2 3.11.5.3.8.7 45.56.11 495 56 ------ 2970 2475 ------- 27720 ; gonna really have to check that ont :)

OpenStudy (amistre64):

27720 .... its good

OpenStudy (anonymous):

your right!

OpenStudy (anonymous):

thank you so much! your my savior!

OpenStudy (amistre64):

youre welcome :)